OpenStudy (anonymous):

If f(x)=2x-4 then find f^-1(x)

OpenStudy (chriss):

solve the equation for x f(x)=2x-4 y=2x-4 y+4=2x \[\frac{y+4}{2}=x\] switch your x's and y's \[y=\frac{x+4}{2}\] \[f^{-1}(x) =\frac{x+4}{2}\]